OpenStudy (anonymous):

a plumber charges $60 for a service call plus $55 for each hour she works. Write a function rule for the total bill for a plumbing job. what is the total bill for job that takes the plumber 3 hours of work?

OpenStudy (anonymous):

cost = $60 + ($55 * (hours-1))

OpenStudy (anonymous):

cost = $60 + ($55 * (2)) cost = $60 + ($110) cost = $170

OpenStudy (anonymous):

Maybe I read that the wrong way. If it's a $60 base charge, PLUS $55 per hour, then it's: fn → cost = $60 + ($55 * (hours)) So: cost = $60 + ($55 * (hours)) cost = $60 + ($55 * 3) cost = $60 + ($165) cost = $225